OpenStudy (anonymous):

f(x) = (x-5)/(9); g(x) = 8x + 5, find g(f(x))

OpenStudy (imer):

\[g(f(x))=8[\frac{( x-5) }{ 9 }]+5\] and then just simplify.

OpenStudy (imer):

The concept is that g(x) is a function and so is f(x) but f(x) is an input for g(x) so you just replace "x" with another function which is g(f(x))

OpenStudy (anonymous):

I got 10.625 = x now.

OpenStudy (anonymous):

8((x-5/9))+5 -5/8 = (x-5)/(9) -5/8 * 9 = x-5 10.625 = x

OpenStudy (anonymous):

That's completely wrong and i Know it.

OpenStudy (anonymous):

A) g(f(x)) = x - (5/8) B) g(f(x)) = x C) g(f(x)) = 8x + 35 D) g(f(x)) = x + 10

OpenStudy (anonymous):

Those are the options.

OpenStudy (imer):

"8((x-5/9))+5 -5/8 = (x-5)/(9) -5/8 * 9 = x-5 10.625 = x" this calculation is wrong.

OpenStudy (anonymous):

What do I have to do?

OpenStudy (imer):

well the answer is not in the options, you sure the question is correct, no typo of sign?

OpenStudy (imer):

its (x-5)/8 not 9 :P

OpenStudy (imer):

the answer is "B"

OpenStudy (anonymous):

Oops, haha thank you. No wonder I was having so much difficulty...

OpenStudy (anonymous):

Even I didn't spot that mistake.

OpenStudy (imer):

\[8[\frac{ (x-5) }{ 8 }]+5 = (x-5)+5=x\] "8" in numerator and denominator get cancelled.

OpenStudy (anonymous):

Thank you.